# Hyperparameter Optimization in Machine Learning

### v1

@@ 2.1.1 Logistic regression for sentiment analysis
As a concrete example, let us revisit in more detail the problem of sentiment analysis introduced in Chapter 1 and, in particular, the logistic regression model considered by yogotama2015-emnlp15. Let us denote the data by $\mathscr{D}:=\bigcup_{N\in\mathbb{N}^{+}}(\mathcal{X}\times\mathcal{Y})^{N}$, where $\mathcal{X}$ is the space of documents and $\mathcal{Y}=\{0,1\}$ the space of labels, expressing a negative or a positive sentiment. The hyperparameter space of the learning algorithm is the Cartesian product $\Lambda=\Lambda_{1}\times\dots\times\Lambda_{7}$ of the seven domains reported in the column “Values” of Table 1.1(a). The first four rows are related to a function $\Phi_{\lambda}:\mathcal{X}\to\mathbb{R}^{n_{\lambda}}$, the feature extractor, that maps documents to $n_{\lambda}$-dimensional vector representation. The last three rows are related to regularization and optimization. The model space is defined as follows:
=$\mathcal{H}=\left\{h_{\lambda}:\mathcal{X}\to\mathcal{Y}\,:\,h_{\lambda}(x)=\textbf{1}\{w^{\intercal}\Phi_{\lambda}(x)\geq 0\},\;w\in\mathbb{R}^{n_{\lambda}},\;\lambda\in\Lambda\right\},$ (2.2)
@@
where $\mathbf{1}$ is the indicator function. In other words, $\mathcal{H}$ is the union of the spaces of all the $n_{\lambda}$-dimensional linear classifiers composed with feature extractor $\Phi_{\lambda}$, as $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ vary in their respective spaces $\Lambda_{1}$, $\Lambda_{2}$, $\Lambda_{3}$ and $\Lambda_{4}$.
@@
In the context of a logistic regression model, learning consists in finding the parameter vector $w$ that minimizes the negative log-likelihood of the data under this model. The first four hyperparameters determine in which space this minimization is carried out. For instance, the larger the range of the n-grams considered (controlled by the hyperparameters $\lambda_{1}$ and $\lambda_{2}$), the more input features and, consequently, parameters the model will have. Unigrams alone ($\lambda_{1}=\lambda_{2}=1$) are very likely not sufficient for sentiment analysis as the sentences $x_{1}=$ the restaurant is not busy but the food is tasty and $x_{2}=$ the restaurant is busy but the food is not tasty are indistinguishable since $\phi_{\lambda}(x_{1})=\phi_{\lambda}(x_{2})$. However, using longer n-grams typically requires increasingly larger corpora to avoid overfitting as the number of possible features (sequences of words of length n) grows exponentially with n.
@@
$\mathcal{J}_{\lambda}(w,\mathcal{D})=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}-\log P\left(y|\Phi_{\lambda}(x);w\right)+\lambda_{6}\Omega_{\lambda_{5}}(w),$ (2.3)
@@
$P\left(1|\Phi_{\lambda}(x);w\right)=\frac{1}{1+e^{-w^{\intercal}\Phi_{\lambda}(x)}}.$
Hyperparameter $\lambda_{5}$ exposes a binary choices between two different forms of regularization: the first favors sparse solutions, the second penalizes components of $w$ with comparatively large values (see Figure 2.1 for a geometric interpretation). The hyperparameter $\lambda_{6}$ controls the strength of such regularization; if, for instance, $\lambda_{5}=L^{1}$, one can expect the number of 00’s in the solution vector to be proportional to the value of $\lambda_{6}$.
@@
The actual hyperparameter space depends on the application and how the problem is framed. For instance, the range of the regularization coefficient $\lambda_{6}$ in our example could in principle take any non-negative real value and it does not need to be limited to the interval $[10^{-5},10^{5}]$. Furthermore, one may want to use both types of regularization simultaneously in (2.3) instead of picking only one of the two, or one may consider the usage of also tetra-grams when extracting features. What to include in $\Lambda$ is a design choice that can have important implications in practice. One of the reasons yogotama2015-emnlp15 succeeded to achieve achieve strong predictive performance with a simple linear classifier is that they included in $\Lambda$ hyperparameters related to the feature extractor $\Phi_{\lambda}$, whereas other authors opted for an a priori fixed map $\Phi$ instead of optimizing over it.
@@ 2.2 The hyperparameter optimization problem
The nested structure of this problem involves minimizing a function whose dependence from its decision variables (the hyperparameters) is implicit through the execution of the learning algorithm. Thus, in principle, to evaluate $f$ at a given point $\lambda$, we must run $A(\mathcal{D},\lambda)$ until convergence. Hence, evaluating the response function can be expensive, as many learning algorithms may take hours, or even days to terminate. We will discuss how we can relax this requirement using low-fidelity approximations in Chapter 5. The second major difficulty inherent to most HPO problems is the irregularity of the search space $\Lambda$. For instance, in our sentiment analysis example, the range of $n$-grams considered by the feature extractor $\Phi_{\lambda}$ is a discrete variable, ruling out the application of continuous algorithms out-of-the-box for optimizing $f$.
When $A$ internally minimizes a training objective, such as the function $\mathcal{J}_{\lambda}$ in the case of logistic regression, problem (2.4)–(2.5) becomes an instance of bilevel programming [colson2007overview, dempe2002foundations, dempe2020bilevel] where the algorithm plays the role of the inner problem [franceschi2018bilevel]. We will return on this view in Chapter 7 when we will discuss gradient-based HPO techniques that explicitly leverage this structure. Here, we note only that we cannot collapse the optimization problem into one level (i.e., joint minimization of both model parameters and hyperparameters). For instance, if we were to optimize the objective $\mathcal{J}_{\lambda}$ from Eq. (2.3) jointly with respect to both $w$ and $\lambda$, we would obtain as solutions the smallest values in $\Lambda$, namely $\lambda_{6}=10^{-5}$ and $\lambda_{7}=10^{-5}$, irrespective of the training set considered. This would defeat the purpose of these two hyperparameters which are limiting the model complexity.
@@ 2.2.1 Estimating generalization
Most of the material and techniques for HPO we covered in this monograph may be used regardless of the specific nature and meaning of $\mathscr{M}$ and, hence, of $f$. However, since many hyperparameters control the inductive bias of their respective learning algorithms, the archetypal setting consists in a response function that estimates the generalization of the models outputby the learning algorithm.
@@
Formalizing this requirement satisfactorily may be a daunting task in certain scenarios (e.g. in generative modelling), but a typical setting is what follows. We assume to have access to a task loss :×$\mathcal{L}:\mathcal{H}\times\mathcal{Z}\to\mathds{R}$, where $\mathcal{Z}$ is an appropriate task-dependent data space and define the generalization error of $h_{\lambda}$ as the expected loss:
$\mathcal{E}(h_{\lambda})\doteq\mathbb{E}_{z\sim p(z)}\left[\mathcal{L}(h_{\lambda},z)\right],$ (2.6)
where $p$ is the unknown task distribution from which the training dataset $\mathcal{D}$ was drawn independently. We may then define the validation error as an empirical estimate of (2.6):
$\mathscr{M}(h_{\lambda},\mathcal{V})=\frac{1}{|\mathcal{V}|}\sum_{z\in\mathcal{V}}\mathcal{L}(h_{\lambda},z),$ (2.7)
where $\mathcal{V}=\{z_{j}\}_{j=1}^{N^{\prime}}$ is a set of points drawn i.i.d. from $p$, dubbed validation, held-out or development set. In the sentiment analysis problem, an instance of supervised binary classification, we have $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$. yogotama2015-emnlp15 use the 00-$1$ error (accuracy), namely $\mathcal{L}(h_{\lambda},(x,y))=\mathbf{1}\{h_{\lambda}(x)\neq y\}$.
@@
When estimating generalization of a model, it is essential to compute $\mathscr{M}$ on a set of points that differs from $\mathcal{D}$. In fact, as learning algorithms adapt their hypotheses through $\mathcal{D}$, evaluating $\mathscr{M}(h_{\lambda},\mathcal{D})$ would result in a distorted estimation of (2.6); e.g., underestimation when the model is under-regularized as illustrated in Figure 2.2, left. This is because in order to obtain an unbiased estimate of the expected loss (2.6) one needs to draw points from the task distribution $p$ independently both from each other and from $h_{\lambda}$.
In practice, one may split the entirety of the available data into two disjoint sets $\mathcal{D}$ and $\mathcal{V}$ and use the first as input to $A$ and hold out the second for evaluating $f$. 222 Note that a third test set would be needed to estimate generalization capabilities of the hypothesis obtained after HPO! More sophisticated approaches, such as cross-validation, involve splitting the data into $K$ subsets and define the response function as a mean over $K$ evaluation of $\mathscr{M}$, whereby at each round one subset is held out for validation and the rest of the data is used for training [stone1974cross, kohavi:1995:study_cross]. Cross-validation may reduce the variance of the estimate of (2.6) at the expense of increased computational cost.
@@ 2.3 Hyperparameter optimization in the literature
  • bergstra_random_2012, later extended in [bergstra_making_2013], brought the attention of the community to random search as a sensible baseline compared to manual or grid search which, at the time of writing, where widespread approaches to HPO. We cover grid, random and quasi-random search techniques in Chapter 3.
  • snoek_practical_2012 used Gaussian-process based Bayesian optimization to optimize the hyperparameters of a 3-layer convolutional neural network [krizhevsky-nips12] on CIFAR-10. They improved the manual found hyperparameter by a human expert which marked state-of-the-art at this point by reducing the test error by $3\%$. This papers sparked interest in Bayesian optimization for HPO in the literature. Chapter 4 is devoted to model-based HPO and Bayesian optimization in the HPO context.
@@
  • jamieson2016non and later li2017hyperband introduced bandit-based approaches to hyperparameter optimization. These methods draw inspiration from the multi-armed bandit literature, particularly the successive halving algorithm [karnin-icml13]. They adaptively allocate resources to promising configurations while terminating unpromising ones early, emphasizing the concepts of budget and efficient compute allocation. These works have been particularly influential in the development of multi-fidelity optimization methods, which we explore in Chapter 5.
@@ 2.4 Practical desiderata for HPO algorithms
The nested optimization formulation (2.4) abstracts away some aspects which are practically relevant and, together, they set aside HPO from other optimization scenarios. For this reason one should weigh an HPO algorithm alongside multiple directions. We summarize these desiderata in Table 2.1.
@@
Table 2.1: List of desirable properties of an HPO algorithm. [We will provide a comparison table of the various HPO techniques where these characteristics will appear as columns after the chapters reviewing the various methods. We will add a link to that table here]
@@
The response function $f(\lambda)$ is typically noisy, depending on random effects in the computation of $A(\mathcal{D},\lambda)$ (e.g., weight initialization, data ordering, e.g., when performing stochastic gradient descent, etc.). It is also computationally demanding to evaluate. Moreover, HPO is typically only a step in a wider iterative process of model development and refinement. This means that in practice, the emphasis in HPO is to achieve a sufficient decrease of $f(\lambda)$ compared to a reference point (e.g., previously used default configuration) within a given budget $b$ – expressed in time, number of function evaluations, monetary costs, etc. – rather than achieving an (approximate) optimum, which is the typical aim in other branches of optimization. Such a decrease needs to be traded off with compute and runtime costs. A good HPO algorithm should be sample efficient, such that it significantly decreases $f$ with as few evaluations of $f$ as possible. More importantly, it should be wall-clock time efficient, meaning that it finds good solutions as rapidly as possible. While number of evaluations and wall-clock time are related, they are not equivalent. For example, decision-making (i.e., choosing the next configuration to evaluate) counts towards wall-clock time and needs to be balanced against evaluation times of $f$ for model-based HPO algorithms. When comparing different algorithms, time efficiency is captured by plotting the best value of $f$ attained against wall-clock time and computing the area under this curve. Some HPO methods tend to witness rapid initial descent of $f$, but may level off at a sub-optimal value, while others spend more of their initial budget in order to ultimately achieve near-optimal performance.
@@
A powerful principle to increase time efficiency is to consider low-fidelity approximations of the response, which are cheaper to compute. For example, if we train a neural network for 64 epochs, we can access $A_{b}(\mathcal{D},\lambda)$, $b<64$, for free. Multi-fidelity HPO methods operate on such a sequence of fidelities $f_{b}(\lambda)$, where $f(\lambda)=f_{64}(\lambda)$. By controlling the budget $b$ for each trial evaluation $\lambda$ (e.g., stopping training after $b$ epochs), they can often decrease $f(\lambda)$ with a much smaller number of full-fidelity evaluations. These methods will be discussed in detail in Chapter 5.

### v2

@@ 2.1.1 Logistic regression for sentiment analysis: an HPO perspective
As a concrete example, let us revisit in more detail the problem of sentiment analysis introduced in Chapter 1 and, in particular, the logistic regression model considered by yogotama2015-emnlp15. Let us denote the data by $\mathscr{D}:=\bigcup_{N\in\mathbb{N}^{+}}(\mathcal{X}\times\mathcal{Y})^{N}$, where $\mathcal{X}$ is the space of documents and $\mathcal{Y}=\{0,1\}$ is the space of labels, expressing a negative or a positive sentiment. The hyperparameter space of the learning algorithm is the Cartesian product $\Lambda=\Lambda_{1}\times\dots\times\Lambda_{7}$ of the seven domains reported in the column “Values” of Table 1.1(a). The first four rows are related to a function $\Phi_{\lambda}:\mathcal{X}\to\mathbb{R}^{n_{\lambda}}$, the feature extractor, that maps documents to $n_{\lambda}$-dimensional vector representation. The last three rows are related to regularization and optimization. The model space is defined as follows:
=$\mathcal{H}=\left\{h_{\lambda}:\mathcal{X}\to\mathcal{Y}\,:\,h_{\lambda}(x)=\textbf{1}_{\mathbb{R}^{+}}(w^{\intercal}\Phi_{\lambda}(x)),\;w\in\mathbb{R}^{n_{\lambda}},\;\lambda\in\Lambda\right\},$ (2.2)
@@
where $\mathbf{1}_{S}$ is the indicator function: $\mathbf{1}_{S}(y)=1$ if $y\in S$ and 00 otherwise. In other words, $\mathcal{H}$ is the union of the spaces of all the $n_{\lambda}$-dimensional linear classifiers composed with feature extractor $\Phi_{\lambda}$, as $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ vary in their respective spaces $\Lambda_{1}$, $\Lambda_{2}$, $\Lambda_{3}$ and $\Lambda_{4}$.
@@
In the context of a logistic regression model, learning consists in finding the parameter vector $w$ that minimizes the negative log-likelihood of the data under this model. The first four hyperparameters determine in which space this minimization is carried out. For instance, the larger the range of the $n$-grams considered, the more input features and, consequently, parameters the model will have. This range is controlled by the hyperparameters $\lambda_{1}$ and $\lambda_{2}$. Unigrams alone ($\lambda_{1}=\lambda_{2}=1$) are very likely not sufficient for sentiment analysis as the sentences $x_{1}=$ the restaurant is not busy but the food is tasty and $x_{2}=$ the restaurant is busy but the food is not tasty are indistinguishable since $\phi_{\lambda}(x_{1})=\phi_{\lambda}(x_{2})$. However, using longer $n$-grams typically requires increasingly larger corpora to avoid overfitting, since the number of possible features (sequences of words of length $n$) grows exponentially with $n$.
@@
$\mathcal{J}_{\lambda}(w,\mathcal{D})=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}-\log\mathbb{P}\left(y|\Phi_{\lambda}(x);w\right)+\lambda_{6}\Omega_{\lambda_{5}}(w),$ (2.3)
@@
$\mathbb{P}\left(1|\Phi_{\lambda}(x);w\right)=\frac{1}{1+e^{-w^{\intercal}\Phi_{\lambda}(x)}}.$
Hyperparameter $\lambda_{5}$ exposes a binary choices between two different forms of regularization: the first favors sparse solutions, the second penalizes components of $w$ with comparatively large values. Figure 2.1 offers a geometric interpretation of these regularization schemes. The hyperparameter $\lambda_{6}$ controls the strength of such regularization; if, for instance, $\lambda_{5}=L^{1}$, one can expect the number of 00’s in the solution vector to be proportional to the value of $\lambda_{6}$.
@@
The actual hyperparameter space depends on the application and how the problem is framed. For instance, the range of the regularization coefficient $\lambda_{6}$ in our example could in principle take any non-negative real value: it does not need to be limited to the interval $[10^{-5},10^{5}]$. Furthermore, one may want to use both types of regularization simultaneously in (2.3) instead of picking only one of the two, or one may consider the usage of also tetra-grams when extracting features. What to include in $\Lambda$ is a design choice that can have important implications in practice. One of the reasons yogotama2015-emnlp15 succeeded to achieve achieve strong predictive performance with a simple linear classifier is that they included in $\Lambda$ hyperparameters related to the feature extractor $\Phi_{\lambda}$, whereas other authors opted for an a priori fixed map $\Phi$ instead of optimizing over it.
@@ 2.2 The hyperparameter optimization problem
The nested structure of this problem involves minimizing a function whose dependence from its decision variables (the hyperparameters) is implicit through the execution of the learning algorithm. Thus, in principle, to evaluate $f$ at a given point $\lambda$, we must run $A(\mathcal{D},\lambda)$ until convergence. Hence, evaluating the response function can be expensive, as many learning algorithms may take hours, or even days to terminate. We will discuss how we can relax this requirement using low-fidelity approximations in Chapter 5. The second major difficulty inherent to most HPO problems is the irregularity of the search space $\Lambda$. For instance, in our sentiment analysis example, the range of $n$-grams considered by the feature extractor $\Phi_{\lambda}$ is a discrete variable, ruling out the application of out-of-the-box continuous algorithms for optimizing $f$.
When $A$ internally minimizes a training objective, such as the function $\mathcal{J}_{\lambda}$ in the case of logistic regression, problem (2.4)–(2.5) becomes an instance of bilevel programming [colson2007overview, dempe2002foundations, dempe2020bilevel] where the algorithm plays the role of the inner problem [franceschi2018bilevel]. We will return on this view in Chapter 7 where we discuss gradient-based HPO techniques that explicitly leverage this structure. A key observation is that we cannot collapse the optimization problem into one level, where we jointly minimize both model parameters and hyperparameters. For instance, if we were to optimize the objective $\mathcal{J}_{\lambda}$ from Eq. (2.3) jointly with respect to both $w$ and $\lambda$, we would obtain as solutions the smallest values in $\Lambda$, namely $\lambda_{6}=10^{-5}$ and $\lambda_{7}=10^{-5}$, irrespective of the training set considered. This would defeat the purpose of these two hyperparameters which are limiting the model complexity.
@@ 2.2.1 Estimating generalization
Most of the material and techniques for HPO we covered in this monograph may be used regardless of the specific nature and meaning of $\mathscr{M}$ and, hence, of $f$. However, since many hyperparameters control the inductive bias of their respective learning algorithms, the archetypal setting consists in a response function that estimates the generalization of the learning algorithm’s output models.
@@
Formalizing this requirement satisfactorily may be a daunting task in certain scenarios (e.g., in generative modelling), but it is easy in the case of supervised learning, where we assume to have access to a task loss :×$\mathcal{L}:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to\mathds{R}$. We can define the generalization error of $h_{\lambda}$ as the expected loss:
$\mathcal{E}(h_{\lambda})\doteq\mathbb{E}_{(x,y)\sim p}\left[\mathcal{L}(h_{\lambda},(x,y))\right],$ (2.6)
where $p$ is the unknown task distribution from which the training dataset $\mathcal{D}$ was drawn independently. Widely used metrics such as classification error or mean-squared-error for regression can be interpreted as empirical estimates of (2.6):
$\mathscr{M}(h_{\lambda},\mathcal{V})=\frac{1}{|\mathcal{V}|}\sum_{(x,y)\in\mathcal{V}}\mathcal{L}(h_{\lambda},(x,y)),$ (2.7)
where $\mathcal{V}$ is a set of points drawn i.i.d. from $p$, dubbed validation, held-out or development set. For example, the sentiment analysis problem addressed in [yogotama2015-emnlp15] is an instance of supervised binary classification and the task loss is the 00-$1$ error, namely $\mathcal{L}(h_{\lambda},(x,y))=\mathbf{1}\{h_{\lambda}(x)\neq y\}$. Note that in many cases, the internal structure of the learning algorithm, $A(\mathcal{D},\lambda)$, also requires the minimization of a loss function $\mathcal{L}^{\prime}$ with respect to the model’s parameters, which may or may not coincide [^2] For example, in the case of classification, the 0-1 loss is intractable to minimize even for linear classifiers and we usually resort to a convex surrogate [nguyen:2009:surrogate_loss_functions] such as the cross-entropy. with $\mathcal{L}$.
@@
When estimating generalization of a model, it is essential to compute $\mathscr{M}$ on a set of points that differs from $\mathcal{D}$. In fact, as learning algorithms adapt their hypotheses through $\mathcal{D}$, evaluating $\mathscr{M}(h_{\lambda},\mathcal{D})$ would result in a distorted estimation of (2.6); e.g., underestimation when the model is under-regularized. This is illustrated in Figure 2.2, left. In order to obtain an unbiased estimate of the expected loss (2.6) one needs to draw points from the task distribution $p$ independently both from each other and from $h_{\lambda}$. Otherwise, one carries over unwanted dependencies between training and validation.
In practice, one may split the entirety of the available data into two disjoint sets $\mathcal{D}$ and $\mathcal{V}$ and use the first as input to $A$ and hold out the second for evaluating $f$. 333 Note that a third test set would be needed to estimate generalization capabilities of the hypothesis obtained after HPO! More sophisticated approaches, such as cross-validation, involve splitting the data into $K$ subsets and define the response function as a mean over $K$ evaluation of $\mathscr{M}$, whereby at each round one subset is held out for validation and the rest of the data is used for training [stone1974cross, kohavi:1995:study_cross]. Cross-validation may reduce the variance of the estimate of (2.6) at the expense of increased computational cost.
@@ 2.2.2 Other types of response functions
2.2.3 General algorithm configuration
Several algorithms, not only those employed in machine learning, require configuration constants that may affect both the quality of the solutions and the computational resources used to find solutions on specific instances [schede:2022:survey_methods_automated]. While in HPO the emphasis is on the former aspect, in general algorithm configuration (AC) the focus is more typically on compute. A typical problem suitable for AC is propositional satisfiability (SAT), where proper choices of parameters can improve solving time significantly [hutter:2009:paramils_automatic_algorithm, hutter_sequential_2011]. Additional interesting examples can be found in the area of discrete optimization, where several parameters may be used to configure metaheuristics [Birattari2009TuningM]. Entering into details would be out of scope for this survey. Still, we would like to point out a striking similarity between AC and HPO. In both cases, optimal (hyper) parameters depend on some unknown probability distribution, such as the data distribution for HPO, or the distribution of future instances of a certain problem in AC. This requires resorting to data (such as a validation set or a set of problem instances) to formulate a proxy empirical optimization problem.
@@ 2.3 Hyperparameter optimization in the literature
  • bergstra_random_2012, later extended in [bergstra_making_2013], brought the attention of the community to random search as a sensible baseline compared to manual or grid search which, at the time of writing, where widespread approaches to HPO. We cover grid, random and more advanced quasi-random search techniques in Chapter 3.
  • snoek_practical_2012 used Gaussian-process based Bayesian optimization to optimize the hyperparameters of a 3-layer convolutional neural network [krizhevsky-nips12] on CIFAR-10. They improved the manual found hyperparameter by a human expert which marked state-of-the-art at this point by reducing the test error by $3\%$. This papers sparked interest in Bayesian optimization for HPO. Chapter 4 is devoted to model-based HPO and Bayesian optimization in the HPO context.
@@
  • jamieson2016non and later li2017hyperband introduced bandit-based approaches to hyperparameter optimization. These methods draw inspiration from the multi-armed bandit literature, particularly the successive halving algorithm [karnin-icml13]. They adaptively allocate resources to promising configurations while terminating unpromising ones early, emphasizing the concepts of budget and efficient compute allocation. These works have been particularly influential in the development of multi-fidelity optimization methods, which we discuss in Chapter 5.
@@ 2.4 Practical desiderata for HPO algorithms
The nested optimization formulation (2.4) abstracts away some aspects which are practically relevant. Together, these characteristics set aside hyperparameter from other optimization scenarios. For this reason one should weigh an HPO algorithm alongside multiple directions. We summarize these desiderata in Table 2.1.
@@
Table 2.1: List of desirable properties of an HPO algorithm.
@@
The response function $f(\lambda)$ is typically stochastic, depending on random effects in the computation of $A(\mathcal{D},\lambda)$ (e.g., weight initialization, data ordering, e.g., when performing stochastic gradient descent, etc.). It is also computationally demanding to evaluate. Moreover, HPO is typically only a step in a wider iterative process of model development and refinement. This means that in practice, the emphasis in HPO is to achieve a sufficient decrease of $f(\lambda)$ compared to a reference point (e.g., previously used default configuration) within a given budget $b$ – expressed in time, number of function evaluations, monetary costs, etc. – rather than achieving an (approximate) optimum, which is the typical aim in other branches of optimization. Such a decrease needs to be traded off with compute and runtime costs. A good HPO algorithm should be sample efficient, such that it significantly decreases $f$ with as few evaluations of $f$ as possible. More importantly, it should be wall-clock time efficient, meaning that it finds good solutions as rapidly as possible. While number of evaluations and wall-clock time are related, they are not equivalent. For example, decision-making (i.e., choosing the next configuration to evaluate) counts towards wall-clock time and needs to be balanced against evaluation times of $f$ for model-based HPO algorithms. When comparing different algorithms, time efficiency is captured by plotting the best value of $f$ attained against wall-clock time and computing the area under this curve. Some HPO methods tend to witness rapid initial descent of $f$, but may level off at a sub-optimal value, while others spend more of their initial budget in order to ultimately achieve near-optimal performance.
@@
A powerful principle to increase time efficiency is to consider low-fidelity approximations of the response, which are cheaper to compute. For example, if we train a neural network for 64 epochs, we can access $A_{b}(\mathcal{D},\lambda)$, that is, the model after $b$ epochs of training have elapsed, for free for all $b\leq 64$. Multi-fidelity HPO methods operate on such a sequence of fidelities $f_{b}(\lambda)$, where, in this case, $f(\lambda)=f_{64}(\lambda)$. By controlling the budget $b$ for each trial evaluation $\lambda$ (e.g., stopping training after $b$ epochs), they can often decrease $f(\lambda)$ with a much smaller number of full-fidelity evaluations. These methods will be discussed in detail in Chapter 5. One aspect that is worth noting when choosing an HPO algorithm is if it can leverage low-fidelity evaluation, and how the quality of such evaluation may impact the final result.